You are titrating 100.0 mL of 0.0400 M Fe2+ in 1 M HCIO4 with 0.100 M Ce4+ to give Fe3+ and Ce3+ using Pt and calomel electro(E) E= 0.767 – 0.05916 log -0.241 (d) Calculate the values of E for the cell when the following volumes of the Ce4+ titrant hNumber 45.0 mL Number 80.0 mL
Answer #1

(a) Balanced titration reaction: Fe2+ + Ce4+ ---> Fe3+ +Ce3+---------------(1)

(b) Half reactions: Ce4+ + e- ---> Ce3+ E0=1.70 V     -----------(1.1)

                            Fe3+ + e- ---> Fe2+   E0=0.767 V -----------(1.2)

(c) Nernst equation for the reaction:

               A++B->A+B+.

Ecell=(E0A+/A-E0B+/B) - 0.05916 log[[B+][A]/[B][A+]] ----------(2)

Here E0A+/A > E0B+/B

and E0B+/B is taken as 0.241 V.

So using Eq. (2), one gets

Ecell=0.767-0.241 - 0.05916 log[[Fe2+]/[Fe3+]]

Ecell=1.70-0.241 - 0.05916 log[[Ce3+]/[Ce4+]].

So the correct answers are (A) and (G).

(d) Before equivalance point:

Nernst equation for the half-reaction (Eq. 1.2):

Ecell=0.767 - 0.05916 log[[Fe2+]/[Fe3+]] -----------(3)

Suppose, V ml of Ce4+ is added.

Moles of Fe2+ present= 100*0.04=4 mM.

After reaction with V*0.1 mM Ce4+, [Fe2+]=4-V*0.1 mM

Number of Fe3+ produced=V*0.1 mM

So from Eq. (3), one has

E=0.767 - 0.05916 log[(4-V*0.1)/(V*0.1)] ----------- (4)

Using Eq. (4), the following result can be obtained:

V=1 ml         E= 0.673 V

V=20 ml      E= 0.767 V

V=38 ml      E= 0.843 V

-------------------------------------------------

V=40 ml      equivalence point

                   E=(1.7+0.767)/2=1.23 V

-------------------------------------

After equivalance point:

Ecell=1.7 - 0.05916 log[[Ce3+]/[Ce4+]] -----------(3)

Initial moles of Fe2+=moles of Ce3+ formed=4 mM

moles of Ce4+=V*0.1- 4.0

Ecell=1.7 - 0.05916 log[4/(V*0.1- 4.0)] -----------(3)

V=45 ml      E= 1.65 V

V=80 ml      E= 1.70 V